how do you calculate the standard deviation and the population deviation for the following You will take a sample of 27 of them for different weeks and you observe the following numbers of hours that they work during a week.numbers? 73,19,16,64,28,28,31,90,60,56,31,56,22,18,45,48,17,17,17,91,92,63,50,51,69,16,17
Answer #1

To find standard deviation we need to find sample mean first

\overline{x}=\frac{\sum x}{n}=43.8889

Create the following table.

data data-mean (data - mean)2
73 29.1111 847.45614321
19 -24.8889 619.45734321
16 -27.8889 777.79074321
64 20.1111 404.45634321
28 -15.8889 252.45714321
28 -15.8889 252.45714321
31 -12.8889 166.12374321
90 46.1111 2126.23354321
60 16.1111 259.56754321
56 12.1111 146.67874321
31 -12.8889 166.12374321
56 12.1111 146.67874321
22 -21.8889 479.12394321
18 -25.8889 670.23514321
45 1.1111 1.23454321
48 4.1111 16.90114321
17 -26.8889 723.01294321
17 -26.8889 723.01294321
17 -26.8889 723.01294321
91 47.1111 2219.45574321
92 48.1111 2314.67794321
63 19.1111 365.23414321
50 6.1111 37.34554321
51 7.1111 50.56774321
69 25.1111 630.56734321
16 -27.8889 777.79074321
17 -26.8889 723.01294321

Find the sum of numbers in the last column to get. (xi – 7)2 = 16620.6667

So Population standard deviation in Σκι - )2 σαν την 16620.6667 - = 24.8109 27

And sample standard deviation is 8 = Σ(εί – )2 16620.6667 η -1 EV 97 -= 25.2835 V